one violin playing at certain location the measured sound pressure level is 50dB. what will be the pressure if four violoins are playing
1
Expert's answer
2020-08-02T15:04:15-0400
"P=10\lg{\frac{I}{I_0}}\\P'=10\lg{\frac{4I}{I_0}}=10\lg{\frac{I}{I_0}}+10\lg{4}\\P'=50+10\lg{4}=56\ dB"